The specific heat of a gas is found to be 0.075 calories at constant volume and its formula wt is 40. The atomicity of the gas would be
 

Grade:Select Grade

2 Answers

Sunil Kumar FP
9 years ago
Sp heat of gas at constant volume is 0.075 cal/g/degC ie0.075cal is absorbed by 1g of it for 1 degree rise in temperature at constant volume.

So total heat absorbed by 1g-mole of the gas is 40*0.075=3 cal.
Since , Cp-Cv=R or Cp=5 so for the gas cp/cv=5/3 and hence the gas is monatomic.
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Sp heat of gas at constant volume is 0.075 cal/g/degC ie0.075cal is absorbed by 1g of it for 1 degree rise in temperature at constant volume.
So total heat absorbed by 1g-mole of the gas is 40*0.075=3 cal.
Since , Cp-Cv=R or Cp=5 so for the gas cp/cv=5/3 and hence the gas is monatomic.
 
Thanks and regards,
Kushagra